OpenStudy (anonymous):

WALK ME THROUGH STEPS Determine the zeros of the quadratic function f(x)=x^2+5x-6. Factor & Zero Product Property

OpenStudy (anonymous):

did you factor it?

OpenStudy (anonymous):

i dont know how ?

OpenStudy (anonymous):

can you find the factors of 6...

OpenStudy (anonymous):

1-6 and 3-2

OpenStudy (anonymous):

yup good now express the coefficient of x which is 5 here in terms of the factors u found..

OpenStudy (anonymous):

I got lost

OpenStudy (anonymous):

why...

OpenStudy (anonymous):

now express the coefficient of x

OpenStudy (anonymous):

you found the factors of 6 .. like 1,6 & 3,2

OpenStudy (anonymous):

yea then

OpenStudy (anonymous):

now 5= 3+2 5= 6-1 if we use 5=3+2 then (3*2=6 , but the constant term free from x is -6) use 5 =6-1 as (here 6* -1 = -6) so we can rewrite f(x)=x^2+5x-6 as f(x)=x^2+(6-1)x-6 f(x) =x^2 +6x -x -6... i hope you can follow...

OpenStudy (anonymous):

I got lost again how did you get f(x) = x^2 +6x -x -6

OpenStudy (anonymous):

given f(x)=x^2+5x-6 so f(x)=x^2+(6-1)x-6 [ noe 5 i have expressed as (6-1) ] now simplify you may have then x^2 +6x -x -6

OpenStudy (anonymous):

now try to take out common terms.. f(x) = x^2 +6x -x -6 f(x) =x(x+6) -1(x+6) again (x+6) is common f(x)= (x+6)(x-1)

OpenStudy (anonymous):

thank you doesn't it bring back to

OpenStudy (anonymous):

yup it should bring back but you need not simplify it again but you have to find factors....